# Stability of discrete-symmetry flocks: sandwich state, traveling domains and motility-induced pinning

### v1

@@ Abstract
$\star$\$$\star$\$footnotetext: These authors have contributed equally to this work
@@ I Introduction
Active matter, a natural or synthetic non-equilibrium system, consists of agents that consume energy to move or exert mechanical forces. Over the past three decades, extensive investigations have established active matter as a prominent area of research Marchetti et al. (2013); De Magistris and Marenduzzo (2015); Bechinger et al. (2016); Shaebani et al. (2020); Bowick et al. (2022); Dauchot (2024); T. Vrugt and Wittkowski (2025). Collections of active particles in such systems often display intricate dynamics and collective phenomena, including the formation of large, ordered structures known as flocks. Flocking Toner , the coordinated movement of numerous agents, is a hallmark of many biological and physical systems. It represents a quintessential non-equilibrium behavior commonly encountered in natural settings Ballerini et al. (2008); Cavagna and Giardina (2014); Becco et al. (2006); Calovi et al. (2014); Steager et al. (2008); Gómez-Nava et al. (2022).
@@
The field of active matter physics arguably originated in 1995, when Vicsek and collaborators published their foundational work Vicsek et al. (1995) on collective motion, drawing an analogy between flocking behavior and ferromagnetism. Their active XY model, later referred to as the Vicsek model (VM), displays true long-range order in two dimensions Toner (2012); Chaté (2020), seemingly in contradiction with the well-known Mermin-Wagner theorem Mermin and Wagner (1966). This does not pose a real inconsistency, as the moving flock represents a far-from-equilibrium system Ginelli (2016); Toner . Later, the active Ising model (AIM) Solon and Tailleur (2013, 2015); Scandolo et al. (2023); Bandyopadhyay et al. (2024) was introduced, which replaces the continuous rotational symmetry of the VM with a discrete symmetry, although preserves the essential physics of the VM. The AIM exhibits three steady-states similar to the VM Solon and Tailleur (2015): disordered gas at high noise and low densities, polar liquid at low noise and high densities, and a phase-separated liquid-gas coexistence region at intermediate densities and noise. The key distinction between the VM and the AIM is that the former exhibits giant density fluctuations leading to microphase separation of the coexistence region, while the latter shows normal density fluctuations resulting in bulk phase separation Solon et al. (2015). Recently, inspired by the AIM, other active spin models such as the $q$-state active Potts model (APM) Chatterjee et al. (2020); Mangeat et al. (2020) and the $q$-state active clock model (ACM) Chatterjee et al. (2022); Solon et al. (2022) have emerged as more generalized models of flocking, which can bridge the continuous symmetry VM and the discrete symmetry AIM Karmakar et al. (2024). Both the APM and ACM are excellent frameworks for investigating collective motion due to their simplicity and capacity to capture key features of flocking behavior, including band-to-lane reorientation transition Chatterjee et al. (2020); Mangeat et al. (2020), flocking to jamming transition under volume exclusion Karmakar et al. (2023), and the role of spatial anisotropy on flocking pattern formation Chatterjee et al. (2022); Solon et al. (2022). All these agent-based models have played an important role in shaping our understanding of polar active matter.
@@
Because of their intrinsically out-of-equilibrium character, active systems are capable of exhibiting long-range order (LRO) and can survive spin wave fluctuations Grégoire and Chaté (2004); Bertin et al. (2006); Chaté et al. (2008a, b); Bertin et al. (2009). This led to the general belief that polar ordered phases in active matter are typically stable against such fluctuations. However, more recent findings have shown that even minor disturbances, such as a small obstacle, can destabilize a flock, triggering transient dynamics before the system reaches a new steady state Codina et al. (2022); Benvegnen et al. (2023); Karmakar et al. (2024). In a similar vein, significant spontaneous fluctuations have also been observed to disrupt the LRO phase. Notable examples include constant-density Toner-Tu flocks Besse et al. (2022), one-species Woo and Noh (2024) and two-species Mangeat et al. (2025) AIM, and the non-reciprocal XY model Dopierala et al. (2025), in which the ordered states eventually become metastable and give way to a disordered phase over long periods – either as irregularly scattered local domains Woo and Noh (2024); Mangeat et al. (2025) or as dynamic aster foams separated by shock lines Besse et al. (2022); Dopierala et al. (2025).
@@
Furthermore, the ordered phase of the VM, which breaks a continuous symmetry, has been argued to be metastable in parts of the phase diagram Codina et al. (2022), where, below a critical noise level, no dense cluster, regardless of its size, triggers a reversal. In contrast, the ordered phase of the AIM, which breaks a discrete symmetry, exhibits metastability both with artificially nucleated minority-phase droplets Benvegnen et al. (2023) and through the spontaneous formation of droplets with opposite polarity Woo and Noh (2024); Mangeat et al. (2025). As a result, the entire flocking phase of the AIM has been proposed to be metastable in the thermodynamic limit Benvegnen et al. (2023). Notably, no spontaneous nucleation or destabilization of the Vicsek liquid phase has been observed so far. This evidence suggests that, unlike in equilibrium systems, flocking models with continuous symmetries may be more robust than their discrete counterparts in non-equilibrium settings Benvegnen et al. (2023). These results raise an important question: Is the stability of two-dimensional orientationally ordered phases in discrete flocking models inherently fragile? If so, under what conditions do these ordered states become stable, and what microscopic mechanisms are responsible for the transition? To explore these questions, it is necessary to investigate a discrete flocking model with more internal degrees of freedom than the AIM. A natural candidate is the more general $q=4$-state APM Chatterjee et al. (2020); Mangeat et al. (2020), which includes four spin states in contrast to the two in the AIM.
@@
In this paper, we investigate the stability of the ordered flocking phase in the 4-state APM with artificial droplet excitations Benvegnen et al. (2023) and the potential for spontaneous nucleation Woo and Noh (2024); Mangeat et al. (2025). Through detailed numerical simulations, we investigate how different parameters, such as droplet size, droplet density, self-propulsion velocity, diffusion constant, and system size, influence the stability of the polar liquid phase. Based on our findings, we argue that polar flocks in the APM are also metastable across a broad region of the phase diagram, consistent with observations in other discrete flocking models Benvegnen et al. (2023); Woo and Noh (2024); Mangeat et al. (2025).
@@ II Model
We consider an ensemble of $N$ particles defined on a two-dimensional lattice of size $L_{x}\times L_{y}$ with periodic boundary conditions. The average particle density of the system is $\rho_{0}=N/L_{x}L_{y}$. The spin state of the $k$-th particle on lattice site $i$ is denoted by $\sigma^{k}_{i}=\{1,2,3,4\}$ corresponding to the movement directions right, up, left, and down, respectively, while the number of particles in state $\sigma$ on-site $i$ is denoted $n^{\sigma}_{i}$. The local density on-site $i$ is then defined by $\rho_{i}=\sum_{\sigma=1}^{4}n^{\sigma}_{i}$. The Hamiltonian of a 4-state APM is defined as $H=\sum_{i}H_{i}$ decomposed as the sum of local Hamiltonian $H_{i}$  Chatterjee et al. (2020); Mangeat et al. (2020):
@@
A particle at site $i$ with state $\sigma$ can either flip to another state $\sigma^{\prime}$ or hop to any neighboring sites. Flipping is a purely on-site phenomenon, and from Eq. (1), one can calculate the local energy difference before and after the flipping. The expression of the energy difference reads $\Delta H_{i}=4J(n_{i}^{\sigma}-n_{i}^{\sigma^{\prime}}-1)/\rho_{i}$. The flipping is then accepted with the rate Chatterjee et al. (2020); Mangeat et al. (2020)
@@
The biased diffusion mechanism is similar to the process described in Refs. Chatterjee et al. (2020); Mangeat et al. (2020). A particle with state $\sigma$ hops to a direction $p$ with rate
@@
The simulation evolves in the unit of Monte Carlo steps (MCS) $\Delta t=[4D+\exp(4\beta)]^{-1}$, resulting from a microscopic time $\Delta t/N$, previously used in the simulations of AIM Solon and Tailleur (2015); Bandyopadhyay et al. (2024); Mangeat et al. (2025), APM Chatterjee et al. (2020); Mangeat et al. (2020), and ACM Chatterjee et al. (2022). During $\Delta t/N,$ a randomly chosen particle either updates its spin state with probability $p_{\rm flip}=W_{\rm flip}\Delta t$ or hops to one of the neighboring sites with probability $p_{\rm hop}=W_{\rm hop}\Delta t$.
@@ III Results
The APM Chatterjee et al. (2020); Mangeat et al. (2020), proposed as a $q$-state extension of the two-state AIM Solon and Tailleur (2013, 2015), displays a flocking transition from a high-temperature, low-density gaseous phase to a low-temperature, high-density polar liquid phase, passing through a liquid-gas coexistence regime at intermediate densities and temperatures. Within this coexistence region, the fully phase-separated liquid bands exhibit a notable reorientation transition—from transverse motion relative to the average direction of particle movement at low velocities and high temperatures, to longitudinal lane alignment at high velocities and low temperatures. This reorientation is driven by the suppression of transverse diffusion at high velocities, which favors the stability of longitudinal lanes. Such a transition is not observed in the AIM Solon and Tailleur (2013, 2015) or the VM Solon et al. (2015), where the bands consistently move in the transverse direction.
@@ III.1 Hydrodynamic theory for droplet-induced perturbations
Figure 1: (color online) Hydrodynamic theory of the APM with droplet excitations. (a–b) Time-evolution snapshots showing (a) the formation of a sandwich state ($\beta=0.9$, $\epsilon=2.4$), and (b) the complete reversal ($\beta=0.75$, $\epsilon=1.5$) for a counter-propagating droplet ($\sigma=1$). (c–d) Time-evolution snapshots showing (c) the complete reversal ($\beta=0.9$, $\epsilon=2.4$), and (d) the formation of a persistent motion with two orthogonal states ($\beta=0.9$, $\epsilon=1.5$) caused by a transversely-propagating droplet ($\sigma=2$). The initial ordered state is of state $\sigma=3$. Parameters of the droplet: $r_{d}=10$ and $\rho_{0}^{d}=5\rho_{0}$. Arrows indicate the direction of motion, and the colorbar represents the corresponding droplet magnetization. A movie (movie01) of the same can be found at Ref. Chatterjee et al. (2025). (e) Magnetization profiles of the sandwich state formed by the counter-propagating droplet ($r_{d}=10$ and $\rho_{0}^{d}=5\rho_{0}$), $\beta=0.9$ and different values of $\epsilon$. (f) Corresponding sandwich width $w_{\rm sw}$ as a function of $\epsilon$. The red, green, and cyan stars represent the corresponding widths of the profiles shown in (c). (g–h) $(\rho_{0}^{d},r_{d})$ stability diagrams for (g) counter- and (h) transversely-propagating droplets for $\epsilon=1.2$. Parameters: $D=1$, $\rho_{0}=3$, and $L=200$.
We first present our results of the droplet excitation from the refined mean-field coarse-grained hydrodynamic description of the APM presented in Ref. Mangeat et al. (2020). If we define the average density of particles in the state $\sigma$ at the 2d position x as $\rho({\bf x},t)=\langle n_{i}^{\sigma}(t)\rangle$, the hydrodynamic equation of the density field $\rho$ for state $\sigma$ is given by Mangeat et al. (2020):
@@
$D_{\parallel}=D(1+\epsilon/3)$ and $D_{\perp}=D(1-\epsilon/3)$ are the diffusion constants in the parallel, ${\bf e_{\parallel}}=(\cos\phi,\sin\phi)$, and perpendicular, ${\bf e_{\perp}}=(\sin\phi,-\cos\phi)$, directions, where $\phi=\pi(\sigma-1)/2$ is the favored angle for state $\sigma$. The self-propulsion velocity $v=4D\epsilon/3$ is along ${\bf e_{\parallel}}$. Derivatives in these directions are denoted as $\partial_{\parallel}={\bf e_{\parallel}}\cdot\nabla$ and $\partial_{\perp}={\bf e_{\perp}}\cdot\nabla$. In the flipping term $I_{\sigma\sigma^{\prime}}$, $\alpha=8(\beta J)^{2}(1-2\beta J/3)$ Mangeat et al. (2020).
@@
We use explicit Euler forward time centered space (FTCS) Press differencing scheme to numerically integrate Eqs. (5). We solve the system of four coupled partial differential equations, one for each $\sigma$, on a square domain of size $L\times L$ with periodic boundary conditions applied in both directions. In our simulation, $L=200$ and the maximum simulation time is $t_{\rm sim}=10^{4}$. To maintain the numerical stability criteria, we set $\Delta x=0.5$ and $\Delta t=10^{-3}/\Delta x^{2}$ as the discretization in space and time, respectively. These discretization parameters satisfy the Courant-Friedrichs-Lewy stability condition Courant et al. (1928). In our numerical implementation, we fix $D=J=r=1$, defining a scaling for the time, temperature, and density, and prepare the initial system with a high-density droplet centered and polarized differently from the surrounding homogeneous liquid phase.
@@
Fig. 1(a–d) shows the droplet dynamics for counter- and transversely-propagating droplets at the low temperature polar ordered phase of the APM. At low temperature ($\beta=0.9$) and high velocity ($\epsilon=2.4$), as $D_{\perp}<D_{\parallel}$, the growth of the droplet along the transverse direction is suppressed, which leads to the formation of a partially reversed mixed sandwich state for a counter-propagating droplet [Fig. 1(a), $t=500$], where regions of the initial ordered phase (blue) coexist with a lane (red) formed by the droplet propagation. However, at higher temperature ($\beta=0.75$) and lower velocity ($\epsilon=1.5$), the counter-propagating droplet grows with a comet-like structure, similar to the one observed for the AIM Benvegnen et al. (2023), and the initial ordered phase is completely reversed [Fig. 1(b), $t=1500$]. Similarly, a complete reversal of the initial liquid phase is observed for a transversely-propagating droplet excitation [Fig. 1(c), $t=10^{4}$]. Moreover, at large $\epsilon$, the pronounced anisotropy in hopping rates, with $D_{\perp}\ll D_{\parallel}$, facilitates the formation of stable longitudinal lanes. However, the expansion of such lanes across the domain is significantly hindered by the suppressed transverse diffusion $D_{\perp}$, resulting in slow lateral growth. Consequently, the steady state for the droplet of $\sigma=2$ in panel (c) is reached much later than that for the droplet of $\sigma=1$ in panel (a). Furthermore, at intermediate velocity ($\epsilon=1.5$), the steady state is composed of two orthogonally moving clusters, with the directions corresponding to the initial ordered state and the inserted droplet. These two clusters never merge into a single flock, even if they frequently collide, due to the square geometry of the periodic domain.
@@
Figure 2: (color online) Time evolution and steady-states with droplet excitations. (i–ii) Schematic illustrating the movement of droplets: (i) a counter-propagating droplet and (ii) a transversely-propagating droplet, relative to the motion of the homogeneous liquid phase. (a, c, e, g, i) Early-time snapshots and (b, d, f, h, j) steady-state snapshots for (a–f) counter-propagating and (g–j) transversely propagating droplets in the 4-state APM. A counter-propagating droplet with a very small self-propulsion velocity of $\epsilon=0.3$ is unable to reverse the initial liquid phase (a–b) but reverses it for intermediate velocity $\epsilon=0.9$ (c–d). (e–f) Counter-propagating droplet at large self-propulsion velocity $\epsilon=2.7$ creates a sandwich state of $\sigma=3$ (blue) and $\sigma=1$ (red). (g–h) Transversely propagating droplet at small self-propulsion velocity $\epsilon=0.9$ could not reverse the initial liquid phase, although reverses it for large self-propulsion velocity $\epsilon=2.7$ (i–j). Parameters: $D=1$, $\beta=1$, $\rho_{0}=10$, $r_{d}=10$, $\rho_{0}^{d}=1.2\rho_{0}$, $L_{x}=500$, and $L_{y}=50$. Colorbar legend: red $(\sigma=1)$: right; green $(\sigma=2)$: up; blue $(\sigma=3)$: left; black $(\sigma=4)$: down. Arrows indicate the direction of motion. A movie (movie02) of the same can be found at Ref. Chatterjee et al. (2025).

### v2

# Stability of discrete-symmetry flocks: Sandwich state, traveling domains, and motility-induced pinning

@@ Abstract
[…]
@@ I Introduction
Active matter, a natural or synthetic non-equilibrium system, consists of agents that consume energy to move or exert mechanical forces. Over the past three decades, extensive investigations have established active matter as a prominent area of research marchetti2013hydrodynamics; de2015introduction; bechinger2016active; shaebani2020computational; bowick2022symmetry; dauchot2024active; te2025metareview. Collections of active particles in such systems often display intricate dynamics and collective phenomena, including the formation of large, ordered structures known as flocks. Flocking toner2024physics, the coordinated movement of numerous agents, is a hallmark of many biological and physical systems. It represents a quintessential non-equilibrium behavior commonly encountered in natural settings ballerini2008interaction; cavagna2014bird; becco2006experimental; calovi2014swarming; steager2008dynamics; gomez2022intermittent.
@@
The field of active matter physics arguably originated in 1995, when Vicsek and collaborators published their foundational work vicsek1995novel on collective motion, drawing an analogy between flocking behavior and ferromagnetism. Their active XY model, later referred to as the Vicsek model (VM), displays true long-range order in two dimensions toner2012reanalysis; chate2020dry, seemingly in contradiction with the well-known Mermin-Wagner theorem mw1966. This does not pose a real inconsistency, as the moving flock represents a far-from-equilibrium system ginelli2016physics; toner2024physics. Later, the active Ising model (AIM) solon2013revisiting; solon2015flocking; AIM2023; AIM2024 was introduced, which replaces the continuous rotational symmetry of the VM with a discrete symmetry, although preserves the essential physics of the VM. The AIM exhibits three steady-states similar to the VM solon2015flocking: disordered gas at high noise and low densities, polar liquid at low noise and high densities, and a phase-separated liquid-gas coexistence region at intermediate densities and noise. The key distinction between the VM and the AIM is that the former exhibits giant density fluctuations leading to microphase separation of the coexistence region, while the latter shows normal density fluctuations resulting in bulk phase separation solon2015phase. Recently, inspired by the AIM, other active spin models such as the $q$-state active Potts model (APM) chatterjee2020flocking; mangeat2020flocking and the $q$-state active clock model (ACM) chatterjee2022polar; solon2022susceptibility have emerged as more generalized models of flocking, which can bridge the continuous symmetry VM and the discrete symmetry AIM karmakar2024consequence. Both the APM and ACM are excellent frameworks for investigating collective motion due to their simplicity and capacity to capture key features of flocking behavior, including band-to-lane reorientation transition chatterjee2020flocking; mangeat2020flocking, flocking to jamming transition under volume exclusion karmakar2023jamming, and the role of spatial anisotropy on flocking pattern formation chatterjee2022polar; solon2022susceptibility. All these agent-based models have played an important role in shaping our understanding of polar active matter.
@@
Because of their intrinsically out-of-equilibrium character, active systems are capable of exhibiting long-range order (LRO) and can survive spin wave fluctuations gregoire2004onset; bertin2006boltzmann; chate2008modeling; chate2008collective; bertin2009hydrodynamic. This led to the general belief that polar ordered phases in active matter are typically stable against such fluctuations. However, more recent findings have shown that even minor disturbances, such as a small obstacle, can destabilize a flock, triggering transient dynamics before the system reaches a new steady state codina2022small; benvegnen2023meta; karmakar2024consequence. In a similar vein, significant spontaneous fluctuations have also been observed to disrupt the LRO phase. Notable examples include constant-density Toner-Tu flocks besse2022metastability, one-species jd2024MIP and two-species TSAIM AIM, and the non-reciprocal XY model dopierala2025inescapable, in which the ordered states eventually become metastable and give way to a disordered phase over long periods – either as irregularly scattered local domains jd2024MIP; TSAIM or as dynamic aster foams separated by shock lines besse2022metastability; dopierala2025inescapable.
@@
Furthermore, the ordered phase of the VM, which breaks a continuous symmetry, has been argued to be metastable in parts of the phase diagram codina2022small, where, below a critical noise level, no dense cluster, regardless of its size, triggers a reversal. In contrast, the ordered phase of the AIM, which breaks a discrete symmetry, exhibits metastability both with artificially nucleated minority-phase droplets benvegnen2023meta and through the spontaneous formation of droplets with opposite polarity jd2024MIP; TSAIM. As a result, the entire flocking phase of the AIM has been proposed to be metastable in the thermodynamic limit benvegnen2023meta. Notably, no spontaneous nucleation or destabilization of the Vicsek liquid phase has been observed so far. This evidence suggests that, unlike in equilibrium systems, flocking models with continuous symmetries may be more robust than their discrete counterparts in non-equilibrium settings benvegnen2023meta. These results raise an important question: Is the stability of two-dimensional orientationally ordered phases in discrete flocking models inherently fragile? If so, under what conditions do these ordered states become stable, and what microscopic mechanisms are responsible for the transition? To explore these questions, it is necessary to investigate a discrete flocking model with more internal degrees of freedom than the AIM. A natural candidate is the more general $q=4$-state APM chatterjee2020flocking; mangeat2020flocking, which includes four spin states in contrast to the two in the AIM.
@@
In this paper, we investigate the stability of the ordered flocking phase in the 4-state APM with artificial droplet excitations benvegnen2023meta and the potential for spontaneous nucleation jd2024MIP; TSAIM. Through detailed numerical simulations, we investigate how different parameters, such as droplet size, droplet density, self-propulsion velocity, diffusion constant, and system size, influence the stability of the polar liquid phase. Based on our findings, we argue that polar flocks in the APM are also metastable across a broad region of the phase diagram, consistent with observations in other discrete flocking models benvegnen2023meta; jd2024MIP; TSAIM.
@@ II Model
We consider an ensemble of $N$ particles defined on a two-dimensional lattice of size $L_{x}\times L_{y}$ with periodic boundary conditions. The average particle density of the system is $\rho_{0}=N/L_{x}L_{y}$. The spin state of the $k$-th particle on lattice site $i$ is denoted by $\sigma^{k}_{i}=\{1,2,3,4\}$ corresponding to the movement directions right, up, left, and down, respectively, while the number of particles in state $\sigma$ on-site $i$ is denoted $n^{\sigma}_{i}$. The local density on-site $i$ is then defined by $\rho_{i}=\sum_{\sigma=1}^{4}n^{\sigma}_{i}$. The Hamiltonian of a 4-state APM is defined as $H=\sum_{i}H_{i}$ decomposed as the sum of local Hamiltonian $H_{i}$  chatterjee2020flocking; mangeat2020flocking:
@@
A particle at site $i$ with state $\sigma$ can either flip to another state $\sigma^{\prime}$ or hop to any neighboring sites. Flipping is a purely on-site phenomenon, and from Eq. (1), one can calculate the local energy difference before and after the flipping. The expression of the energy difference reads $\Delta H_{i}=4J(n_{i}^{\sigma}-n_{i}^{\sigma^{\prime}}-1)/\rho_{i}$. The flipping is then accepted with the rate chatterjee2020flocking; mangeat2020flocking
@@
The biased diffusion mechanism is similar to the process described in Refs. chatterjee2020flocking; mangeat2020flocking. A particle with state $\sigma$ hops to a direction $p$ with rate
@@
The simulation evolves in the unit of Monte Carlo steps (MCS) $\Delta t=[4D+\exp(4\beta)]^{-1}$, resulting from a microscopic time $\Delta t/N$, previously used in the simulations of AIM solon2015flocking; AIM2024; TSAIM, APM chatterjee2020flocking; mangeat2020flocking, and ACM chatterjee2022polar. During $\Delta t/N,$ a randomly chosen particle either updates its spin state with probability $p_{\rm flip}=W_{\rm flip}\Delta t$ or hops to one of the neighboring sites with probability $p_{\rm hop}=W_{\rm hop}\Delta t$.
@@ III Results
The APM chatterjee2020flocking; mangeat2020flocking, proposed as a $q$-state extension of the two-state AIM solon2013revisiting; solon2015flocking, displays a flocking transition from a high-temperature, low-density gaseous phase to a low-temperature, high-density polar liquid phase, passing through a liquid-gas coexistence regime at intermediate densities and temperatures. Within this coexistence region, the fully phase-separated liquid bands exhibit a notable reorientation transition—from transverse motion relative to the average direction of particle movement at low velocities and high temperatures, to longitudinal lane alignment at high velocities and low temperatures. This reorientation is driven by the suppression of transverse diffusion at high velocities, which favors the stability of longitudinal lanes. Such a transition is not observed in the AIM solon2013revisiting; solon2015flocking or the VM solon2015phase, where the bands consistently move in the transverse direction.
@@ III.1 Hydrodynamic theory for droplet-induced perturbations
Figure 1: (color online) Hydrodynamic theory of the APM with droplet excitations. (a–b) Time-evolution snapshots showing (a) the formation of a sandwich state ($\beta=0.9$, $\epsilon=2.4$), and (b) the complete reversal ($\beta=0.75$, $\epsilon=1.5$) for a counter-propagating droplet ($\sigma=1$). (c–d) Time-evolution snapshots showing (c) the complete reversal ($\beta=0.9$, $\epsilon=2.4$), and (d) the formation of a persistent motion with two orthogonal states ($\beta=0.9$, $\epsilon=1.5$) caused by a transversely-propagating droplet ($\sigma=2$). The initial ordered state is of state $\sigma=3$. Parameters of the droplet: $r_{d}=10$ and $\rho_{0}^{d}=5\rho_{0}$. Arrows indicate the direction of motion, and the colorbar represents the corresponding droplet magnetization. A movie (movie01) of the same can be found at Ref. zenodo. (e) Magnetization profiles of the sandwich state formed by the counter-propagating droplet ($r_{d}=10$ and $\rho_{0}^{d}=5\rho_{0}$), $\beta=0.9$ and different values of $\epsilon$. (f) Corresponding sandwich width $w_{\rm sw}$ as a function of $\epsilon$. The red, green, and cyan stars represent the corresponding widths of the profiles shown in (c). (g–h) $(\rho_{0}^{d},r_{d})$ stability diagrams for (g) counter- and (h) transversely-propagating droplets for $\epsilon=1.2$. Parameters: $D=1$, $\rho_{0}=3$, and $L=200$.
We first present our results of the droplet excitation from the refined mean-field coarse-grained hydrodynamic description of the APM presented in Ref. mangeat2020flocking. If we define the average density of particles in the state $\sigma$ at the 2d position x as $\rho({\bf x},t)=\langle n_{i}^{\sigma}(t)\rangle$, the hydrodynamic equation of the density field $\rho$ for state $\sigma$ is given by mangeat2020flocking:
@@
$D_{\parallel}=D(1+\epsilon/3)$ and $D_{\perp}=D(1-\epsilon/3)$ are the diffusion constants in the parallel, ${\bf e_{\parallel}}=(\cos\phi,\sin\phi)$, and perpendicular, ${\bf e_{\perp}}=(\sin\phi,-\cos\phi)$, directions, where $\phi=\pi(\sigma-1)/2$ is the favored angle for state $\sigma$. The self-propulsion velocity $v=4D\epsilon/3$ is along $\bf e_{\parallel}$}. Derivatives in these directions are denoted as $\partial_{\parallel}={\bf e_{\parallel}}\cdot\nabla$ and $\partial_{\perp}={\bf e_{\perp}}\cdot\nabla$. In the flipping term $I_{\sigma\sigma^{\prime}}$, $\alpha=8(\beta J)^{2}(1-2\beta J/3)$ mangeat2020flocking.
@@
We use explicit Euler forward time centered space (FTCS) press2007numerical differencing scheme to numerically integrate Eqs. (5). We solve the system of four coupled partial differential equations, one for each $\sigma$, on a square domain of size $L\times L$ with periodic boundary conditions applied in both directions. In our simulation, $L=200$ and the maximum simulation time is $t_{\rm sim}=10^{4}$. To maintain the numerical stability criteria, we set $\Delta x=0.5$ and $\Delta t=10^{-3}/\Delta x^{2}$ as the discretization in space and time, respectively. These discretization parameters satisfy the Courant-Friedrichs-Lewy stability condition courant1928partiellen. In our numerical implementation, we fix $D=J=r=1$, defining a scaling for the time, temperature, and density, and prepare the initial system with a high-density droplet centered and polarized differently from the surrounding homogeneous liquid phase.
@@
Fig. 1(a–d) shows the droplet dynamics for counter- and transversely-propagating droplets at the low temperature polar ordered phase of the APM. At low temperature ($\beta=0.9$) and high velocity ($\epsilon=2.4$), as $D_{\perp}<D_{\parallel}$, the growth of the droplet along the transverse direction is suppressed, which leads to the formation of a partially reversed mixed sandwich state for a counter-propagating droplet [Fig. 1(a), $t=500$], where regions of the initial ordered phase (blue) coexist with a lane (red) formed by the droplet propagation. However, at higher temperature ($\beta=0.75$) and lower velocity ($\epsilon=1.5$), the counter-propagating droplet grows with a comet-like structure, similar to the one observed for the AIM benvegnen2023meta, and the initial ordered phase is completely reversed [Fig. 1(b), $t=1500$]. Similarly, a complete reversal of the initial liquid phase is observed for a transversely-propagating droplet excitation [Fig. 1(c), $t=10^{4}$]. Moreover, at large $\epsilon$, the pronounced anisotropy in hopping rates, with $D_{\perp}\ll D_{\parallel}$, facilitates the formation of stable longitudinal lanes. However, the expansion of such lanes across the domain is significantly hindered by the suppressed transverse diffusion $D_{\perp}$, resulting in slow lateral growth. Consequently, the steady state for the droplet of $\sigma=2$ in panel (c) is reached much later than that for the droplet of $\sigma=1$ in panel (a). Furthermore, at intermediate velocity ($\epsilon=1.5$), the steady state is composed of two orthogonally moving clusters, with the directions corresponding to the initial ordered state and the inserted droplet. These two clusters never merge into a single flock, even if they frequently collide, due to the square geometry of the periodic domain.
@@
Figure 2: (color online) Time evolution and steady-states with droplet excitations. (i–ii) Schematic illustrating the movement of droplets: (i) a counter-propagating droplet and ($i$) a transversely-propagating droplet, relative to the motion of the homogeneous liquid phase. (a, c, e, g, i) Early-time snapshots and (b, d, f, h, j) steady-state snapshots for (a–f) counter-propagating and (g–j) transversely propagating droplets in the 4-state APM. A counter-propagating droplet with a very small self-propulsion velocity of $\epsilon=0.3$ is unable to reverse the initial liquid phase (a–b) but reverses it for intermediate velocity $\epsilon=0.9$ (c–d). (e–f) Counter-propagating droplet at large self-propulsion velocity $\epsilon=2.7$ creates a sandwich state of $\sigma=3$ (blue) and $\sigma=1$ (red). (g–h) Transversely propagating droplet at small self-propulsion velocity $\epsilon=0.9$ could not reverse the initial liquid phase, although reverses it for large self-propulsion velocity $\epsilon=2.7$ (i–j). Parameters: $D=1$, $\beta=1$, $\rho_{0}=10$, $r_{d}=10$, $\rho_{0}^{d}=1.2\rho_{0}$, $L_{x}=500$, and $L_{y}=50$. Colorbar legend: red $(\sigma=1)$: right; green $(\sigma=2)$: up; blue $(\sigma=3)$: left; black $(\sigma=4)$: down. Arrows indicate the direction of motion. A movie (movie02) of the same can be found at Ref. zenodo.